1: You need to make an aqueous solution of 0.212 M barium hydroxide for an experiment in lab, using a 300 mL volumetric flask. How much solid barium hydroxide should you add? ____grams

2: How many milliliters of an aqueous solution of 0.219 M chromium(III) sulfate is needed to obtain 3.70 grams of the salt?____ml

3: In the laboratory you dissolve 23.1 g of zinc nitrate in a volumetric flask and add water to a total volume of 250 . mL. What is the molarity of the solution? ______ M
Answer #1

1)

volume , V = 3*10^2 mL

= 0.3 L

use:

number of mol,

n = Molarity * Volume

= 0.212*0.3

= 6.36*10^-2 mol

Molar mass of Ba(OH)2,

MM = 1*MM(Ba) + 2*MM(O) + 2*MM(H)

= 1*137.3 + 2*16.0 + 2*1.008

= 171.316 g/mol

use:

mass of Ba(OH)2,

m = number of mol * molar mass

= 6.36*10^-2 mol * 1.713*10^2 g/mol

= 10.9 g

Answer: 10.9 g

Only 1 question at a time please